### v1

@@ 2 A network description of code
2.1 Nomenclature and definitions
The goal of computer programs is to implement algorithms on a computer. An algorithm is a terminating sequence of operations which takes an input and computes an output using memory to record interim results. [^10] Memory to record interim instructions is sometimes called a “scratch pad”, in line with early definitions of algorithms which pre-date computers. See, for example, Knuth (1997) (Ch.1). We use the term broadly to include algorithms that rely heavily on user inputs and are highly interactive (e.g. websites, spreadsheet and text processing software, servers). To implement an algorithm, programming languages need to provide a means of reading input and writing output, have a list of instructions that can be executed by a computer, and provide a means of storing values in memory. More formally:
@@ 2.1 Nomenclature and definitions
Software developers, which we refer to as coders, use programming languages to implement algorithms. Early programs were written in programming languages like FORTRAN and later in C, both of which adhere to the procedural programming paradigm. In this paradigm, code is developed via procedures that can communicate with each other via calls and returns.
Definition 2.
A procedure is a sequence of programming instructions, which make use of the resources provided by the computer system, to perform a specific task.
Procedures are an integral tool of almost all programming languages because they allow a logical separation of tasks and abstraction from low-level instructions, which greatly reduces the complexity of writing code. We use the term procedure broadly and explicitly allow procedures to return values. The first large software systems like the UNIX, Linux, and Windows operating systems or the Apache web server have been developed using procedural programming.
Today, however, the dominant modern software development paradigm is Object-oriented programming (OOP). [^12] For a principal discussion of object-oriented programming and some differences to procedural programming, see Abelson et al. (1996). Kay (1993) provides an excellent historical account of the development of early object-oriented programming. Under this paradigm, code is developed primarily in classes. Classes contain data in the form of variables and data structures as well as code in the form of procedures (which are often called methods in the OOP paradigm). Classes can communicate with one another via procedures using calls and returns.
@@
Access to a computer’s memory is managed in most programming languages through the use of variables, which are memory locations that can hold a value. A variable has a scope which describes where in a program text a variable can be used. [^13] There are three different types of variables in object-oriented code. First, a member variable can take a different value for each object of a class. Second, a class variable has the same value for all objects of a class. And third, a global variable has the same value for all objects (i.e. irrespective of the object’s class) in the program. The extent of a variable defines when a variable has a meaningful value during a program’s run-time. Depending on the programming language, variables also have a type, which means that only certain kinds of values can be stored in a variable. [^14] For example, C— is a statically typed programming language, i.e. the C— compiler checks during the compilation of the source code that variables are only passed values for storage that the variable type permits. Similarly, Rust— is a statically typed programming language that, unlike C—, is also object oriented. Python—, on the other hand, is dynamically typed, and checks the validity of value assignments to variables only during run-time.
Similar to variables, data structures are memory locations that hold a collection of data, stored in a way that implements logical relationships among the data. For example, an array is a data structure whose elements can be identified by an index. Different programming languages permit different operations on data structures, for example the appending to and deleting from an array.
@@ 2.2 Dependency Graphs
Most open source software is organized by package managers like Cargo (for the Rust programming language) that standardize and automate software distribution. To make the installation of complex software projects easier for users, package managers also keep track of a package’s dependencies and allow users to install these alongside the package they are installing. The existence of package managers and the level of automation they provide is necessary because modern software is frequently updated and different versions of the same package are not always compatible. Packages are logically grouped into repositories, controlled and managed by a so-called maintainer. [^15] Maintainers can also be organizations. Some repositories are also controlled by a group of maintainers, but for our purposes, this is immaterial.
Our unit of analysis is the network of dependencies among repositories. Specifically, we describe a software system consisting of $\mathcal{N}=\{1,\ldots,N\}$ repositories with $N=|\mathcal{N}|\geq 3$, each of which is managed by a different maintainer who decides whether to implement code herself or to re-use existing code from other repositories. This (re-)use of existing code gives rise to linkages between repositories. The resulting network $\mathcal{G}=(\mathcal{N},\mathcal{E})$ with $\mathcal{E}\subseteq\mathcal{N}\times\mathcal{N}$ is called the dependency graph of the software system. Denote $g_{ij}=1$ if $(i,j)\subseteq\mathcal{E}$ is an existing link from repo $i$ to repo $j$, indicating that $i$ depends on $j$. Denote as $g=\{g_{ij}\}$ the resulting $N\times N$ adjacency matrix of the repository dependency graph $\mathcal{G}$. We exclude self-loops by defining $g_{ii}\coloneqq 0$, because, technically, a package cannot depend on itself.
@@
There are two reasons for constructing the dependency graph on the repository rather than the package level. First, information about the popularity of software packages is only created on the repository level, not the package level. [^16] For example, users can “star” a repository on GitHub if they find it particularly useful. And second, repositories provide a logical grouping of related software packages. Alternatively, we could construct the dependency graph on the level of individual software packages. However, the breadth of code and functionality between packages can be much smaller than the breadth of code and functionality within a package. Yet another alternative is to study call graphs arising from procedures within the same software package (see, for example, Grove et al. (1997)). But these are state-dependent, i.e., dependencies arise during runtime and depending on how the package is executed and interacted with.
Conveniently, the website libraries.io provides information collected from various open source software package mangers, including Cargo. [^17] They obtain this data by scraping publicly available repositories hosted on GitHub, GitLab, or Bitbucket. The data provided includes information on pro-jects—essentially the same as a Cargo package—as well as repositories and for each repository a list of all dependencies, defined on the project level. Each project belongs to exactly one repository, which allows us to construct dependencies among repositories from the provided data.
@@
On the repository level, this data includes information about the Size of the repository in kB, its Popularity, measured as the number of stars it has on the website hosting the repository, and as an alternative measure of popularity the number of contributors, i.e. the number of individual coders who have added code to a repository, raised or answered an issue, or reviewed code submitted by others. [^18] We restrict ourselves to repositories that have a size larger than zero and that have no more than $50,0000$ stars where the latter restriction is included to eliminate 154 repositories where the number of stars seems to have been artificially and unreasonably increased using e.g. a script.
@@
$35,473$ $696,790$ $91$ $35,274$ $696,679$
Table 1: Network statistics for the dependency graph of the Cargo Rust ecosystem. #Nodes and #Edges is the number of nodes and edges for the total network and for the largest weakly connected component, respectively. #Components is the number of connected components. LCC indicates the largest (weakly) connected components.
In-Degree Out-Degree Eigenvector
Centrality
Mean $19.75$ $19.75$ $0.0013$
Std. Dev. $267.57$ $35.91$ $0.018$
Minimum 00 00 $0.0$
p5 00 00 $0.0$
Median 00 $5$ $0.0$
P95 $8$ $100$ $0.0005$
Maximum $14,585$ $316$ $1.0$
Table 2: Distribution of in- and out-degree and eigenvector centrality for the largest weakly connected component of the repo-based dependency graph of the Cargo Rust ecosystem with $N=35,274$ nodes.
Figure 1: Dependency graph of the Cargo Rust ecosystem. Each node is a repository in the libraries.io dataset, and each edge is a dependency between two package versions, each of which belongs to a different repository. Color indicates different modularity classes identified using Blondel et al. (2008).
Table 1 gives a high-level overview of the repo-based dependency graph constructed in this way. While there are $91$ weakly connected components, the largest weakly connected component covers almost all (over $99\%$) nodes and edges. This means that we will still cover all relevant dynamics even if we focus on the largest weakly connected component of the dependency graph only.
Furthermore, there is a lot of heterogeneity in how interconnected repositories are, as Table 2 shows. The standard deviation of the out-degree, i.e. how many dependencies a given repository uses, is almost twice as large as the mean, meaning that there is substantial heterogeneity, with the median being that a repository uses $5$ dependencies, while the maximum is $316$ dependencies. This heterogeneity is even more stark for the in-degree, i.e. how often a given repository is a dependency for another repository. The standard deviation is more than ten times as large as the mean and the median is 00, while the maximum is $14,585$.
In Table 3 we show the $10$ most depended-on packages (i.e. with the highest in-degree). For each of them it is quite intuitive why the repository is used so often. The most-depended on repository, for example, is libc, which allows Rust code to interoperate with C code, which is particularly important for high-performance and security relevant computations. The second-most depended repository is rand, which is used to generate random numbers, while log is used for run-time logging, a crucial task during code development.
@@
Repo name Eigenvector In-Degree Popularity Size #Contrib.
Centrality
rust-lang/libc 1 14,585 674 14,323 338
rust-random/rand 0.764 11,244 546 5,224 546
rust-lang/log 0.723 10,628 674 1,106 68
retep998/winapi-rs 0.679 10,117 705 19,643 128
bitflags/bitflags 0.686 10,105 248 782 47
serde-rs/serde 0.738 10,072 2,764 6,468 132
r[$\ldots$]y/lazy-static.rs 0.692 10,045 709 920 41
BurntSushi/byteorder 0.531 7,831 464 250 38
serde-rs/json 0.498 7,272 1,293 1,654 69
rust-lang/regex 0.490 7,261 1,291 13,175 116
Table 3: List of the ten most depended-on repositories for the Rust Cargo ecosystem. Popularity is measured as the number of stars a repository received on the website hosting it. Size is measured in Kilobytes of all code in the repository. #Contrib. is the number of coders who have contributed to a repository.
The largest weakly connected component of the repo-based dependency graph is shown in Figure 1, where we apply the community detection algorithm of Blondel et al. (2008) to color-code packages in the same community. The algorithm identifies $32$ distinct communities where community members are relatively highly connected to other community members and less connected to nodes outside the community.
2.3 Covariate Data
In addition to the software dependency data, we also obtain additional data from the from libraries.io website for each package (e.g. https://libraries.io/cargo/libc), which we show for the ten most-depended on repositories in Table 3. We use four types of covariates, summarized in a vector of observable attributes $\bm{x}_{i}$. First, we use a repository’s Size of all code in the repository, measured in Kilobytes. Second, Popularity is measured as the number of stars–an expression of how many people like a repository or find it useful–a repository received on the website hosting it, e.g. github. And, third, we use an alternative popularity measure, Num.Contributors, defined as the number of coders who have contributed to a repository. A contribution is not only writing code, but contributing to the code generation process more broadly. Contributiions are defined as code reviews, code commits, creating issues related to a repository, and pull requests, i.e. proposing changes to a repository.
@@ 2.3 Covariate Data
Size [kB] 3,903.91 28,152.89 1 3 79 9,672 1,025,320
Popularity 30.85 392.90 0 0 0 63 41,652
#Contributors 3.49 17.86 0 0 1 10 1,247
Table 4: Distribution of covariates for the $N=35,274$ repositories in the largest connected component of our sample. Size is measured in Kilobytes of all code in the repository. Popularity is measured as the number of stars a repository received on the website hosting it. #Contributors is the number of coders who have contributed to a repository.
To facilitate the estimation of our model, we create a categorical variable for each of our covariates using quartiles of the distribution. The reason to discretize the variables is mostly computational, as our algorithm is based on stochastic blockmodels with discrete types (Bickel et al., 2013; Vu et al., 2013). If the covariates are discrete, then the whole machinery for estimation of blockmodels can be adapted to estimate the unobserved heterogeneity (Vu et al., 2013; Babkin et al., 2020; Dahbura et al., 2021); while with continuous variables the computational costs of estimation become prohibitive for such large networks. [^19] We provide more details on this in the model and the appendix. More technical details are provided in Babkin et al. (2020) and Dahbura et al. (2021).
@@ 3 Model
We model the decisions of a team of coders contributing to a software repository consisting of closely related software packages. Coders decide whether to use functionality from other packages or develop it themselves. In addition to the four observables introduced above, also assume that there are $K$ discrete types, unobservable to the researchers but observable to other coders, $\bm{z}_{i}=(z_{i1},...,z_{iK})$. The type of a software package can be thought of as the basic function of the package and examples include operating system components, graphical user interfaces, text processing software, compilers, and so on. It can also capture unobservable quality of the code and/or developers. A package of type $k$ is denoted by $z_{ik}=1$ and $z_{i\ell}=0$ for all $\ell\neq k$. We use notation $\bm{x}$ and $\bm{z}$ to denote the matrix of observable and unobservable characteristics for all the packages.
@@
The utility function of coder $i$ from network $\bm{g}$, observables $\bm{x}$, unobservables $\bm{z}$ and parameters $\bm{\theta}=(\bm{\alpha},\bm{\beta},\bm{\gamma})$ is:
$U_{i}\left(\bm{g},\bm{x},\bm{z};\bm{\theta}\right)=\sum_{j=1}^{N}g_{ij}u_{ij}(\alpha,\beta)+\sum_{j=1}^{N}\sum_{r\neq i,j}^{N}g_{ij}g_{jr}w_{ijr}(\gamma)+\sum_{j=1}^{N}\sum_{r\neq i,j}^{N}g_{ij}g_{ri}v_{ijr}(\gamma),$ (1)
The payoff $u_{ij}(\bm{\alpha},\bm{\beta}):=u\left(\bm{x}_{i},\bm{x}_{j},\bm{z}_{i},\bm{z}_{j};\bm{\alpha},\bm{\beta}\right)$ is the direct utility of linking to package $j$. It is a function of observables ($\bm{x}_{i},\bm{x}_{j}$), unobservables ($\bm{z}_{i},\bm{z}_{j}$) and parameters ($\bm{\alpha},\bm{\beta}$). This is the benefit of creating the dependencies to package $j$, net of costs—a coder will have to audit the code of the linked package and determine it’s quality—of maintaining the link. The cost also includes modification and adaptation of the code that the coder has to do to be able to use the functions and methods available in package $j$. Furthermore, using another coder’s code also increases the risk that the code contains an undetected bug (assuming that coders care more about the code of their own package than the code of their dependencies), which can be captured as a cost. And lastly, re-using code, while common practice in modern software development, means coders require fewer skills, which might negatively impact their future productivity and can thus be interpreted as another cost.
We assume that the utility function $u_{ij}(\bm{\alpha},\bm{\beta})$ is parameterized as follows
@@
$u_{ij}(\bm{\alpha},\bm{\beta})=\begin{cases}\alpha_{w}+\displaystyle\sum_{p=1}^{P}\beta_{wp}\bm{1}\{\bm{x}_{ip}=\bm{x}_{jp}\}&\text{ if }z_{i}=z_{j}\\ \alpha_{b}+\displaystyle\sum_{p=1}^{P}\beta_{bp}\bm{1}\{\bm{x}_{ip}=\bm{x}_{jp}\}&\text{otherwise,}\end{cases}$ (2)
where the intercept $\alpha$ is interpreted as the cost of of forming the dependencies, and it is a function of unobservables only; and $\bm{1}\{\bm{x}_{ip}=\bm{x}_{jp}\}$ are indicators functions equal to one when the $p$-th covariates are the same for $i$ and $j$. In this parsimonious specification we allow the net benefits of link to vary with both observed and unobserved heterogeneity.

### v2

@@ 2 A network description of code
2.1 The open source software paradigm
The goal of computer programs is to implement algorithms on a computer. An algorithm is a terminating sequence of operations which takes an input and computes an output using memory to record interim results. [^11] Memory to record interim instructions is sometimes called a “scratch pad”, in line with early definitions of algorithms which pre-date computers. See, for example, Chapter 1 of Knuth (1997). We use the term broadly to include algorithms that rely heavily on user inputs and are highly interactive (e.g. websites, spreadsheet and text processing software, servers). Algorithms are implemented on a computer using code. Formally:
@@ 2.1 The open source software paradigm
Software developers, which we refer to as coders, use programming languages to implement algorithms. Today, the dominant modern software development paradigm is object-oriented programming (OOP). [^13] For a principal discussion of object-oriented programming and some differences to procedural programming, see Abelson et al. (1996). Kay (1993) provides an excellent historical account of the development of early object-oriented programming. Under this paradigm, code is developed primarily in classes, which contain data in the form of variables and data structures as well as code in the form of procedures. [^14] Also called functions, or methods in the OOP paradigm. Classes can communicate with one another via procedures using calls and returns.
@@ 2.2 Dependency Graphs
Most open source software is organized by package managers like PyPi for the Python programming language that standardize and automate software distribution. To make the installation of complex software projects easier for users, package managers keep track of a package’s dependencies and allow users to install these alongside the package they are installing. The existence of package managers and the level of automation they provide is necessary because modern software is frequently updated and different versions of the same package are not always compatible. Packages are logically grouped into repositories, controlled and managed by a maintainer. [^15] Maintainers can also be organizations. Some repositories are also controlled by a group of maintainers, but for our purposes, this is immaterial.
[…]
@@
There are two reasons for constructing the dependency graph on the repository rather than the package level. First, information about the popularity of a piece of software is relevant for maintainers linking decisions, but only created on the repository level. [^17] For example, users can “star” a repository on GitHub if they find it particularly useful. And second, repositories provide a logical grouping of closely related software packages. Alternatively, we could construct the dependency graph on the level of individual software packages. However, the breadth of code and functionality between packages can be much smaller than the breadth of code and functionality within a package. [^18] Yet another alternative is to study call graphs arising from procedures within the same software package (see, for example, Grove et al., 1997). But these are state-dependent, i.e., dependencies arise during runtime and depending on how the package is executed and interacted with. Conveniently, the website libraries.io provides information collected from various open source software package managers, including PyPi. [^19] The website libraries.io obtains this data by scraping publicly available repositories hosted on GitHub, GitLab, and Bitbucket. The data provided includes information on projects–essentially the same as a PyPi package–as well as repositories and for each repository a list of all dependencies, defined on the project level. Since each project belongs to exactly one repository, we construct dependencies among repositories from the data provided.
@@
On the repository level, this data includes information about the size of the repository in kB (Size), its popularity, measured as the number of stars it has on the website hosting the repository (Popularity), and as an alternative measure of popularity the number of contributors, i.e. the number of individual coders who have added code to a repository, raised or answered an issue, or reviewed code submitted by others. [^20] We restrict ourselves to repositories that have a size above the 5% percentile of the size distribution, and that have more than $2$ stars, but no more stars than the 95% percentile of the distribution of stars.
@@
$17,095$ $53,498$ $422$ $16,102$ $52,897$
Table 1: Network statistics for the dependency graph of the PyPi ecosystem.
Mean Std.Dev. Min. p5 Median P95 Max.
In-Degree $3.29$ $50.49$ 00 00 00 $6$ $4,210$
Out-Degree $3.29$ $5.0$ 00 00 $2$ $12$ $95$
Table 2: Distribution of in- and out-degree for the largest weakly connected component of the repo-based dependency graph of the PyPi ecosystem with $N=16,102$ nodes.
Figure 1: Dependency graph of the PyPi ecosystem.
Table LABEL:Table::NetworkStats provides a high-level overview of the repository-based dependency graph constructed in this way. While there are $91$ weakly connected components, the largest weakly connected component covers almost all (over $99\%$) nodes and edges. This means that we still cover all relevant dynamics even by only focusing on the largest weakly connected component of the dependency graph.
Furthermore, Table 2 shows that there is a lot of heterogeneity in how interconnected repositories are. The standard deviation of the out-degree, i.e. how many dependencies a given repository uses, is almost twice as large as the mean. Therefore, we observe substantial heterogeneity among repositories, with the median repository using $2$ dependencies and the maximum being $95$ dependencies. This heterogeneity is even more pronounced for the in-degree, i.e. how often a given repository is a dependency for another repository. The standard deviation is more than ten times as large as the mean; the median is 00, while the maximum is $14,585$.
In Table LABEL:Tab::TopRepos we show the $10$ most depended-on packages (i.e. with the highest in-degree) The repository with highest in-degree is six, which facilitates compatibility between python version 2 and 3 (where some major updates have led to serious incompatibility issues). The second-highest in-degree repository is pytest-cov, which is used to check a code’s test coverage, which is a crucial step in modern software development. Likewise, the third-most commonly linked software package is mock, which is also used for software testing.
@@
Repo name In-Degree Popularity Size Maturity
benjaminp/six 4,210 579 1,749 1,106
pytest-dev/pytest-cov 2,584 643 777 2,097
testing-cabal/mock 2,002 421 1,306 1,728
pypa/wheel 1,572 156 1,302 902
kjd/idna 1,336 106 507 2,422
pytest-dev/pluggy 952 364 471 1,721
pypa/packaging 782 141 540 2,069
pallets/markupsafe 775 282 175 3,492
PyCQA/mccabe 767 242 82 2,518
coveralls-clients/coveralls-python 761 406 328 2,532
Table 3: List of the ten most depended-on repositories for the PyPi ecosystem.
The largest weakly connected component of the repo-based dependency graph is shown in Figure LABEL:Fig::RustDependencyGraph, where we apply the community detection algorithm of Blondel et al. (2008) to color-code packages in the same community. The three largest communities contain about $10-11\%$ of all repositories each.
2.3 Observables: Covariates
In addition to the software dependency data, we obtain additional data from the from libraries.io website for each package (e.g. https://libraries.io/pypi/pandas), which we show for the ten most-depended on repositories in Table LABEL:Tab::TopRepos. We use three types of covariates, summarized in a vector of observable attributes $\bm{x}_{i}$. First, we use a repository’s Size of all code in the repository, measured in Kilobytes. Second, Popularity is measured as the number of stars–an expression of how many people like a repository or find it useful–a repository received on the website hosting it, e.g., GitHub. And, third, the time in days between the first release of a repository and the reference date of 13 January 2020 is denoted Maturity.
@@ 2.3 Observables: Covariates
Size [kB] 7,729.8 52,557.84 6 23 273 27,627 3,484,077
Popularity 70.49 133.35 3 3 16 364 853
Maturity 1384.6 838.14 4 248 1250 2960 4270
Table 4: Distribution of covariates
To facilitate the estimation of our model, we create a categorical variable for each of our covariates using quartiles of the distribution. The reason for discretizing variables is computational, as our algorithm is based on stochastic block models with discrete types (Bickel et al., 2013; Vu et al., 2013). If the covariates are discrete, the estimation engine of stochastic block models can be adapted to estimate the unobserved heterogeneity (Vu et al., 2013; Babkin et al., 2020; Dahbura et al., 2021). Otherwise, when the covariates are continuous, the computational costs are prohibitive for such large networks. [^21] We provide more details in the appendix; see also Babkin et al. (2020) and Dahbura et al. (2021).
@@ 3 Model
We model software development as a dynamic process, assuming that the developers update their packages to add or improve their functionality. Maintainers can update their packages by either adding functionalities to their own packages or by creating dependencies to other packages that have the desired functionalities.
3.1 Setup
Each package $i$ and its developer team is characterized by a vector of $P$ observable attributes $\bm{x}_{i}\in\mathbb{R}^{P}$, which are stored in a matrix $\bm{x}\in\mathbb{R}^{N\times P}$. We assume that each of the $N$ packages belongs to one of $K\geq 2$ types, unknown to researchers but known to other developers. The type of a software package may be its purpose, which may be known to developers but may not be known to researchers, or may correspond to unobservable qualities of the code and its developer team. To indicate the type $k$ of package $i$, we define $z_{ik}\coloneqq 1$ and $z_{il}\coloneqq 0$ for all $l\neq k$ if package $i$ is of type $k$, and write $\bm{z}_{i}\coloneqq(z_{i1},\ldots,z_{iK})$. We assume that the unobservables are independent and identically distributed according to a multinomial distribution:
[…]
where $\bm{\eta}\coloneqq(\eta_{1},\ldots,\eta_{K})$ is a vector of $K$ probabilities summing to $1$. We represent the software dependency network $\bm{g}$ by the adjacency matrix, with elements $g_{ij}=1$ if package $i$ depends on package $j$ and $g_{ij}=0$ otherwise. In general, software dependency networks are directed: If package $i$ depends on package $j$, then package $j$ need not depend on package $i$.
@@ 3.1 Setup
The utility that developers of package $i$ receive from the software dependency network $\bm{g}$, with observables $\bm{x}$, unobservables $\bm{z}$, and parameters $\bm{\theta}\coloneqq(\bm{\alpha},\,\bm{\beta},\,\gamma)$, is of the form
$U_{i}(\bm{g};\,\bm{x},\,\bm{z},\,\bm{\theta})\;\coloneqq\;\sum_{j=1}^{N}g_{ij}u_{ij}(\bm{\alpha},\,\bm{\beta})+\sum_{j=1}^{N}\,\sum_{r\neq i,j}^{N}g_{ij}g_{jr}v_{ijr}(\gamma)+\sum_{j=1}^{N}\,\sum_{r\neq i,j}^{N}g_{ij}g_{ir}w_{ijr}(\gamma).$ (2)
The function $u_{ij}(\bm{\alpha},\,\bm{\beta})\coloneqq u(\bm{x}_{i},\,\bm{x}_{j},\bm{z}_{i},\,\bm{z}_{j},\bm{\alpha},\,\bm{\beta})$ is the direct utility of linking package $i$ to package $j$. It is a function of observables ($\bm{x}_{i},\,\bm{x}_{j}$), unobservables ($\bm{z}_{i},\,\bm{z}_{j}$), and parameters ($\bm{\alpha},\,\bm{\beta}$). This represents the benefit of a dependency of package $i$ on package $j$, net of costs: e.g., the developers will have to audit the code of package $j$ and determine its quality, and have to maintain the dependency on package $j$. The cost may includes modification and adaptation of the code, because the developers have to be able to use the functions and methods available in package $j$. Additional costs may arise from relying on other packages, because other packages may contain undetected bugs. Finally, relying on existing code is convenient but may decrease the incentives of developers to maintain and improve coding skills, which in turn can undermine the future productivity of developers and can hence be viewed as a cost.
Motivated by the application to software dependency networks, we parameterize $u_{ij}$ as follows:
@@
$u(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha},\,\bm{\beta})\;\coloneqq\;\begin{cases}u_{w}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{w},\,\bm{\beta}_{w})\ &\text{ if both $i$ and $j$ are of type $k$}\vspace{.25cm}\\ u_{b}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{b},\,\bm{\beta}_{b})\ &\text{otherwise,}\end{cases}$ (3)
where the utility for dependencies within types $u_{w}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{w},\,\bm{\beta}_{w})$ is
$\begin{array}[]{lllll}u_{w}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{w},\,\bm{\beta}_{w})&\coloneqq&\alpha_{w1}+\alpha_{w1}^{\ell}\,\log(N_{k})+\displaystyle\sum_{p=1}^{P}\beta_{wp}\,h_{ijp}(\bm{x}_{i},\bm{x}_{j})\\ &+&\displaystyle\sum_{p=1}^{P}\beta_{wp}^{\ell}\,h_{ijp}(\bm{x}_{i},\bm{x}_{j})\,\log(N_{k}),\end{array}$ (4)
and the utility for dependencies between types $u_{b}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{b},\,\bm{\beta}_{b})$ is
$\begin{array}[]{lllll}u_{b}(\bm{x}_{i},\bm{x}_{j},\,\bm{z}_{i},\bm{z}_{j},\,\bm{\alpha}_{b},\,\bm{\beta}_{b})&\coloneqq&\alpha_{b}+\displaystyle\sum_{p=1}^{P}\beta_{bp}\,h_{ijp}(\bm{x}_{i},\bm{x}_{j}).\end{array}$ (5)
The terms in the within- and between-type utilities can be interpreted as follows: The intercept $\alpha_{w1}\in\mathbb{R}$ represents the fixed cost of forming a dependency; the term $\alpha_{w2}\,\log(N_{k})$ with weight $\alpha_{w2}\in\mathbb{R}$ is the part of the cost that depends on the size $N_{k}$ of type $k$; [^22] In principle, the specification can be extended to incorporate aggregated terms at the type-level. In this sense, $\log(N_{k})$ is one among other examples of terms that can be used to make the model more flexible. For instance, one may include the average of observed characteristics of type $k$ as one of the covariates. We do not pursue this specification here, given the very sparse nature of the software dependency network. and $h_{ijp}(\bm{x}_{i},\bm{x}_{j})\coloneqq\bm{1}\{{x}_{ip}={x}_{jp}\}$ is an indicator function, which is equal to $1$ if the $p$th covariate of $i$ and $j$ matches and is 00 otherwise. Size-dependent terms of the form $\log(N)$ were suggested by Krivitsky et al. (2011), motivated by invariance considerations (encouraging the expected number of links to be invariant to $N$), and adapted to observed types by Krivitsky et al. (2023) and to unobserved types by Babkin et al. (2020). The micro-behavioral foundations of size-dependent terms of the form $\log(N)$ were studied by Butts (2019). The vectors $\bm{\alpha}\coloneqq(\alpha_{w1},\,\alpha_{w2},\,\alpha_{b})\in\mathbb{R}^{3}$ and $\bm{\beta}\coloneqq(\beta_{w1},\,\ldots,\,\beta_{wP},\,\beta_{b1},\,\ldots,\,\beta_{bP},\beta_{w1}^{\ell},\,\ldots,\,\beta_{wP}^{\ell})\in\mathbb{R}^{3P}$ are parameters to be estimated.